In how many ways can a committee of 4 people be chosen out of 8 people ?
Answer by jim_thompson5910(35256)

There are 8*7*6*5 = 1680 ways to do it if order mattered


However, on a committee, order usually doesn't matter. So you have to divide by 4! = 4*3*2*1 = 24 to get 1680/24 = 70


So the final answer is 70